FZR • Postfach 51 01 19 • 01314 Dresden

FAX

Organisationseinheit

(zweizeilig) Position Name

Telefon (03 51) 2 60-00 00 Telefax (03 51) 2 60-00 00 V.Name@fz-rossendorf.de

16. Januar 2001

To:  An alle Mitarbeiterinnen und Mitarbeiter

FAX-NR.:  (03 51) 2 60-???
Pages to follow: 1

FAX-Text